OpenStudy (anonymous):

Simplify the expression (4x − 3)(x + 5).

OpenStudy (xxdragoonxx):

= 4x*x+4x*5−3*x−3*5 = 4x*x+5*4x−3*x−3*5 = 4x^2+20x−3x−15 = 4x^2+17x-15